# Strategic Informed Trading and the Value of Private Information

### v1

@@ Appendix B Proofs from Section 1 and Appendix A
Derivation of $\widehat{\psi}_{I,\iota}(g,z)$ in Theorem 1.6.
From (10), (11), $V(y)=\widehat{\Pi}$, (19) and (20) we obtain
$\begin{split}\widehat{\psi}_{I,\iota}(g,z)&=-p_{0}+\frac{p_{I}}{1+\widehat{y}}(g-h_{\iota})+\frac{p_{I}}{1+2\widehat{y}}(h_{\iota}-p_{0}),\\ &=-\left(1+\frac{p_{I}}{1+\widehat{y}}\right)p_{0}+\frac{p_{I}}{1+\widehat{y}}g-\frac{p_{I}\widehat{y}}{(1+\widehat{y})(1+2\widehat{y})}(h_{\iota}-p_{0}),\\ &=-\left(1+\frac{p_{I}}{1+\widehat{y}}\right)p_{0}+\frac{p_{I}}{1+\widehat{y}}g-\frac{1+p_{I}}{1+\widehat{y}}(p_{\iota}(h_{\iota})-p_{0}),\\ &=\frac{1}{1+\widehat{y}}\left(p_{I}g-(1+p_{I})p_{\iota}(h_{\iota})-\widehat{y}p_{0}\right).\end{split}$
Above, the third equality used (20).
@@
The law of $X$ given $\sigma(G,Z_{N})$ has density
(65) $\frac{\mathbb{P}\left[X\in dx\big{|}\sigma(G,Z_{N})\right]}{\mathbb{P}\left[X\in dx\right]}=\frac{\mathbb{P}\left[X\in dx\big{|}\sigma(G)\right]}{\mathbb{P}\left[X\in dx\right]}=\left(\frac{e^{-\frac{1}{2}x^{\prime}P_{I}x+x^{\prime}P_{I}g}}{\mathbb{E}\left[e^{-\frac{1}{2}X^{\prime}P_{I}X+X^{\prime}P_{I}g}\right]}\right)\bigg{|}_{g=G}.$
Using this in (47), on $\left\{G=g,Z_{N}=z\right\}$ the insider minimizes over $\psi=\psi(g,z)$ with $\psi(G,Z_{N})\in{\mathcal{A}}_{I}$
$\begin{split}&(\psi-\psi_{I,0})^{\prime}p_{\iota}\left(\psi,z\right)+\log\left(\frac{\mathbb{E}\left[e^{-\psi^{\prime}X-\frac{1}{2}X^{\prime}P_{I}x+X^{\prime}P_{I}g}\right]}{\mathbb{E}\left[e^{-\frac{1}{2}X^{\prime}P_{I}X+X^{\prime}P_{I}g}\right]}\right).\\ \end{split}$
As $X\sim N(\mu_{X},P_{X}^{-1})$, this specifies to
$(\psi-\psi_{I,0})^{\prime}p_{\iota}\left(\psi,z\right)+\frac{1}{2}\psi^{\prime}(P_{I}+P_{X})^{-1}\psi-\psi^{\prime}\left(P_{I}+P_{X}\right)^{-1}\left(P_{I}g+P_{X}\mu_{X}\right).$
Plugging in for $p_{\iota}$, using $\psi^{\prime}M\psi=(1/2)\psi^{\prime}(M+M^{\prime})\psi$, and grouping by powers of $\psi$ gives
(66) $\begin{split}&\frac{1}{2}\psi_{I,0}^{\prime}M\left(\psi_{I,0}-\frac{z}{\alpha_{I}}\right)-\psi_{I,0}^{\prime}V+\frac{1}{2}\psi^{\prime}\left(M+M^{\prime}+(P_{I}+P_{X})^{-1}\right)\psi\\ &\qquad-\psi^{\prime}(P_{I}+P_{X})^{-1}\left(P_{I}g+P_{X}\mu_{X}+(P_{I}+P_{X})\left((M+M^{\prime})\psi_{I,0}-M\frac{z}{\alpha_{I}}-V\right)\right).\end{split}$
Plugging in for $M$ and using $P_{X|G}$ from (48), the optimizer $\widehat{\psi}_{I}$ is
(67) $\begin{split}\widehat{\psi}_{I,\iota}(g,z)&=P_{X|G}^{1/2}({\mathcal{Y}}+{\mathcal{Y}}^{\prime}+1_{d})^{-1}P_{X|G}^{-1/2}\bigg{(}P_{I}g+P_{X}\mu_{X}-P_{X|G}V+P_{X|G}^{1/2}({\mathcal{Y}}+{\mathcal{Y}}^{\prime})P_{X|G}^{-1/2}\psi_{I,0}\\ &\quad-P_{X|G}^{1/2}{\mathcal{Y}}P_{X|G}^{-1/2}\frac{z}{\alpha_{I}}\bigg{)}.\end{split}$
The identity in (49) with ${\mathcal{M}},\Lambda_{\iota},\mathcal{V}$ from (50) follow by direct computations. ∎
@@
As the uninformed is a price taker, this follows immediately from Proposition A.6 below, with the appropriate substitutions $p\to p_{\iota}$, $H\to H_{\iota}$. ∎
@@
Throughout we suppress the dependence of all quantities on ${\mathcal{Y}}$. Now, (49), (52) and (53) show the right side of (54) is affine in the signal $H_{\iota}$, and after dividing by $\alpha_{U}$ and right-multiplying by ${\mathcal{M}}^{-1}$, (56) is precisely the equation which eliminates the $H_{\iota}$ terms. As for the constant terms, we need
@@
$\begin{split}\Pi&=\left(\alpha_{I}1_{d}-\alpha_{U}(P_{U,\iota}+P_{X})M\right){\mathcal{M}}\mathcal{V}+\alpha_{I}\psi_{I,0}+\alpha_{U}\left(P_{X}\mu_{X}-(P_{U,\iota}+P_{X})V\right).\end{split}$
As (56) does not involve $V$, if $\widehat{{\mathcal{Y}}}$ is such that (56) holds, the above becomes
$\begin{split}\Pi&=-\alpha_{U}P_{U,\iota}\mathcal{V}+\alpha_{I}\psi_{I,0}+\alpha_{U}\left(P_{X}\mu_{X}-(P_{U,\iota}+P_{X})V\right),\\ &=-\alpha_{U}P_{I}^{-1}P_{U,\iota}\left(P_{X}\mu_{X}-\psi_{I,0}-P_{X|G}V\right)+\alpha_{I}\psi_{I,0}+\alpha_{U}\left(P_{X}\mu_{X}-(P_{U,\iota}+P_{X})V\right),\\ &=\alpha_{U}(P_{I}-P_{U,\iota})P_{I}^{-1}P_{X}(\mu_{X}-V)+(\alpha_{I}P_{I}+\alpha_{U}P_{U,\iota})\psi_{I,0},\end{split}$
where the last equality follows from (50), (48) and simplifying. Therefore, using $\widehat{\Pi}$ and $p_{0}$ from (55), and $P_{\iota}$ from (57) we deduce
@@
$\begin{split}\widehat{\Pi}&=\frac{\alpha_{U}}{\alpha_{I}+\alpha_{U}}\left(P_{I}-P_{U,\iota}\right)P_{I}^{-1}P_{X}\left(\mu_{X}-V\right)+\frac{\alpha_{U}P_{U,\iota}+\alpha_{I}P_{I}}{\alpha_{I}+\alpha_{U}}P_{I}^{-1}\psi_{I,0},\\ &=\left(P_{I}-P_{\iota}\right)P_{I}^{-1}P_{X}\left(p_{0}-V\right)+\left(P_{I}-P_{\iota}\right)P_{I}^{-1}\widehat{\Pi}+P_{\iota}P_{I}^{-1}\psi_{I,0}.\end{split}$
This yields
$V=p_{0}+P_{X}^{-1}P_{I}\left(P_{I}-P_{\iota}\right)^{-1}P_{\iota}P_{I}^{-1}\left(\psi_{I,0}-\widehat{\Pi}\right).$
(57) now follows as $P_{I}(P_{I}-P_{\iota})^{-1}P_{\iota}P_{I}^{-1}=P_{\iota}(P_{I}-P_{\iota})^{-1}$. As for (58), we have from (52) that
$p_{\iota}(h_{\iota})=p_{0}+M{\mathcal{M}}(h_{\iota}-p_{0})+\left(M{\mathcal{M}}-1_{d}\right)p_{0}+M{\mathcal{M}}\mathcal{V}+V.$
Next,
@@
$\begin{split}&\left(M{\mathcal{M}}-1_{d}\right)p_{0}+M{\mathcal{M}}\mathcal{V}+V\\ &\qquad=\left(M{\mathcal{M}}-1_{d}\right)p_{0}+V+M{\mathcal{M}}P_{I}^{-1}\left(P_{X}p_{0}+\widehat{\Pi}-\psi_{I,0}-P_{X|G}V\right),\\ &\qquad=\left(M{\mathcal{M}}-1_{d}\right)p_{0}+p_{0}+P_{X}^{-1}P_{\iota}(P_{I}-P_{\iota})^{-1}\left(\psi_{I,0}-\widehat{\Pi}\right)\\ &\qquad\qquad+M{\mathcal{M}}P_{I}^{-1}\left(P_{X}p_{0}-\left(\psi_{I,0}-\widehat{\Pi}\right)-P_{X|G}p_{0}-P_{X|G}P_{X}^{-1}P_{\iota}(P_{I}-P_{\iota})^{-1}\left(\psi_{I,0}-\widehat{\Pi}\right)\right).\end{split}$
Because $P_{X|G}=P_{I}+P_{X}$ the $p_{0}$ terms cancel out. The $\psi_{I,0}-\widehat{\Pi}$ terms are
$\begin{split}&-M{\mathcal{M}}P_{I}^{-1}+\left(1_{d}-M{\mathcal{M}}P_{I}^{-1}(P_{X}+P_{I})\right)P_{X}^{-1}P_{\iota}(P_{I}-P_{\iota})^{-1}\\ &\qquad=\left(P_{X}^{-1}-M{\mathcal{M}}(P_{\iota}^{-1}+P_{X}^{-1})\right)P_{\iota}(P_{I}-P_{\iota})^{-1}.\end{split}$
@@
Assume the market signal takes the form $H=G+\Lambda Z_{N}$, for a to-be-determined matrix $\Lambda$, and the time 00 price is a function $p=p(H)$ of the market signal. Clearly, $\sigma(G,H)=\sigma(G,Z_{N})$, and hence (21) is equivalent to
$\begin{split}&\inf_{\psi\in\sigma(G,Z_{N})}\mathbb{E}\left[e^{-\psi_{I,0}^{\prime}p(H)-\psi^{\prime}\left(X-p(H)\right)}\big{|}\sigma(G,Z_{N})\right].\end{split}$
We may ignore the $\psi_{I,0}^{\prime}p(H)$ term, and using (65), on the set $\left\{G,H\right\}=\left\{g,h\right\}$ the insider solves
$\inf_{\psi}\left(\psi^{\prime}p(h)+\log\left(\frac{\mathbb{E}\left[e^{-\psi^{\prime}X-\frac{1}{2}X^{\prime}P_{I}X+X^{\prime}P_{I}g}\right]}{\mathbb{E}\left[e^{-\frac{1}{2}X^{\prime}P_{I}X+X^{\prime}P_{I}g}\right]}\right)\right).$
Using that $X\sim N(\mu_{X},P_{X}^{-1})$, this problem is equivalent to minimizing
$\frac{1}{2}\psi_{I}^{\prime}(P_{I}+P_{X})^{-1}\psi_{I}-\psi_{I}^{\prime}(P_{I}+P_{X})^{-1}\left(P_{X}\mu_{X}+P_{I}g-(P_{I}+P_{X})p(h)\right),$
which yields $\widehat{\psi}_{I}$ from (62). Next, the clearing condition (4) mandates that $\widehat{\psi}_{I}+Z_{N}/\alpha_{I}$ be $\sigma(H)$ measurable in equilibrium. Using (62), this in turn implies
$\widehat{\psi}_{I}(G,H)+\frac{Z_{N}}{\alpha_{I}}=P_{I}\left(G+\frac{1}{\alpha_{I}}P_{I}^{-1}Z_{N}\right)+P_{X}\mu_{X}-(P_{I}+P_{X})p(H),$
must be $\sigma(H)$ measurable, which is achieved setting $H$ as in (60). Turning to $U$, as $H$ is of the same form as $G$, we may repeat the steps for the insider, but using $(P_{U},H)$ rather than $(P_{I},G)$, to conclude that $\widehat{\psi}_{U}$ as in (62) is optimal. We now use (4) to identify $p(H)$. Indeed, in equilibrium we must have
$\Pi=\alpha_{I}\left(\widehat{\psi}_{I}(G,H)+\frac{Z_{N}}{\alpha_{I}}\right)+\alpha_{U}\widehat{\psi}_{U}(H),$
which, using $\widehat{\Pi}$ from (2) and $P$ from (61), implies
$\begin{split}\widehat{\Pi}&=\frac{\alpha_{I}}{\alpha_{I}+\alpha_{U}}\left(P_{I}H+p_{X}\mu_{X}-(P_{I}+P_{X})p(H)\right)+\frac{\alpha_{U}}{\alpha_{I}+\alpha_{U}}\left(P_{U}H+p_{X}\mu_{X}-(P_{U}+P_{X})p(H)\right),\\ &=PH+P_{X}\mu_{X}-(P_{X}+P)p(H).\end{split}$
Therefore, using (2) again, $p(H)=(P_{X}+P)^{-1}\left(PH+P_{X}p_{0}\right)$ which gives (61), finishing the result.
The proof of Proposition A.7, as well as those in both Sections C and D simplify if we adjust the notation in (17) by defining
(68) $\beta:=\kappa p_{I}=\alpha_{I}^{2}p_{I}p_{N};\quad\lambda_{I}:=\lambda=\frac{\alpha_{I}}{\alpha_{I}+\alpha_{U}};\quad\lambda_{U}:=1-\lambda=\frac{\alpha_{U}}{\alpha_{I}+\alpha_{U}},\quad R:=\frac{1}{1+p_{I}}.$
@@
Next, recall that in the general setup of Appendix A
(70) $X=\mu_{X}+P_{X}^{-1/2}\mathcal{E}_{X};\quad G=X+\frac{1}{\sqrt{p_{I}}}P_{X}^{-1/2}\mathcal{E}_{I};\qquad Z_{N}=\frac{1}{\sqrt{p_{N}}}P_{X}^{1/2}\mathcal{E}_{N},$
where $\mathcal{E}_{X},\mathcal{E}_{I},\mathcal{E}_{N}$ are three independent $N(0,1_{d})$ random variables.
@@
We use the notation in (68), and suppress $\widehat{y}$ from the $M,{\mathcal{M}},\mathcal{V},V$ functions. We start with the price taking equilibria. From (45), (60), (68), as well as the general pricing formulas in (73) below we obtain after some computation

### v2

@@ B.1. Proofs from Section 1
B.1.1. Theorem 1.2 from Theorem A.5
Throughout, we use that in Section 1, $P_{X}=1,\mu_{X}=0$. We also use Assumption A.4. We start with the market signal. Here, (9) follows from (59) and (62) (throughout at $y=\widehat{y}$). The precision $p_{U,\iota}$ in (10) is immediate from (7) and (62). The price in (11) is immediate from (60) and (62).
The insider and uninformed trader’s policy in (13) follows from (61) and (8) using (62). The slope $M_{p,\iota}$ follows from (62), and the intercept $V_{p,\iota}$ follows from (45) and (8) given the translation $V_{p,\iota}=V_{0,p,\iota}+p_{{\mathbb{Q}}_{0}}$ and the fact we showed (see right below (56)) that $V_{0,p,\iota}=0$.
B.1.2. Proposition 1.3 from Proposition A.6
Throughout, we again use that in Section 1, $P_{X}=1,\mu_{X}=0$, and we enforce Assumption A.4. The market signal in (15) follows immediately from (65). The precision $p_{U}$ in (16) is immediate from (69). The price $p$ in (17) follows directly from (7), (67) and (69). The trading strategies for $I,U$ are immediate given (68) and (8).
B.1.3. Proposition 1.4 from Proposition A.7
This is an immediate consequence of Proposition A.7 given Assumption A.4 and that in Section 1, $P_{X}=1,\mu_{X}=0$.
B.2. Proofs from Appendix A
To ease notation, throughout all proofs of this section we omit the $p,\iota$ subscript from $M_{p,\iota},V_{0,p,\iota}$ in the price response function $p_{\iota}$ from (48), and we omit the functional dependence on ${\mathcal{Y}}$.
@@ B.2. Proofs from Appendix A
The ${\mathbb{Q}}_{0}$ law of $X_{0}$ given $\sigma(G_{0},Z_{N})$ has density
(71) $\frac{\mathbb{Q}_{0}\left[X_{0}\in dx\big{|}\sigma(G_{0},Z_{N})\right]}{{\mathbb{Q}}_{0}\left[X_{0}\in dx\right]}=\left(\frac{e^{-\frac{1}{2}x^{\prime}P_{I}x+x^{\prime}P_{I}g}}{\mathbb{E}^{{\mathbb{Q}}_{0}}\left[e^{-\frac{1}{2}X_{0}^{\prime}P_{I}X+X_{0}^{\prime}P_{I}g}\right]}\right)\bigg{|}_{g=G_{0}}.$
Using this in (48), on $\left\{G_{0}=g,Z_{N}=z\right\}$ the insider minimizes over $\theta=\theta(g,z)$
$\begin{split}&\theta^{\prime}\left(M\left(\theta+\frac{z}{\alpha_{I}}\right)+V_{0}\right)+\log\left(\frac{\mathbb{E}^{{\mathbb{Q}}_{0}}\left[e^{-\theta^{\prime}X_{0}-\frac{1}{2}X_{0}^{\prime}P_{I}X_{0}+X_{0}^{\prime}P_{I}g}\right]}{\mathbb{E}^{{\mathbb{Q}}_{0}}\left[e^{-\frac{1}{2}X_{0}^{\prime}P_{I}X_{0}+X_{0}^{\prime}P_{I}g}\right]}\right).\\ \end{split}$
As $X_{0}\stackrel{{\scriptstyle{\mathbb{Q}}_{0}}}{{\sim}}N(0,P_{X}^{-1})$, this specifies to (using $\theta^{\prime}M\theta=(1/2)\theta^{\prime}(M+M^{\prime})\theta$)
(72) $\begin{split}&\frac{1}{2}\theta^{\prime}\left(M+M^{\prime}+P_{X|G}^{-1}\right)\theta-\theta^{\prime}P_{X|G}^{-1}\left(P_{I}g-P_{X|G}\left(M\frac{z}{\alpha_{I}}+V_{0}\right)\right).\end{split}$
Using $M$ as in (49), the optimizer $\widehat{\theta}_{I,\iota}$ is
(73) $\begin{split}\widehat{\theta}_{I,\iota}(g,z)&=P_{X|G}^{1/2}({\mathcal{Y}}+{\mathcal{Y}}^{\prime}+1_{d})^{-1}P_{X|G}^{-1/2}\bigg{(}P_{I}g-P_{X|G}V_{0}-P_{X|G}^{1/2}{\mathcal{Y}}P_{X|G}^{-1/2}\frac{z}{\alpha_{I}}\bigg{)}.\end{split}$
The identity in (50) with ${\mathcal{M}},\Lambda_{\iota},\mathcal{V}$ from (51) follow by direct computation. ∎
@@
Using (52) one can show the ${\mathbb{Q}}_{0}$ law of $X_{0}$ given $\sigma(H_{0,\iota})$ has density
(74) $\frac{\mathbb{Q}_{0}\left[X_{0}\in dx\big{|}\sigma(H_{0,\iota})\right]}{{\mathbb{Q}}_{0}\left[X_{0}\in dx\right]}=\left(\frac{e^{-\frac{1}{2}x^{\prime}P_{U,\iota}x+x^{\prime}P_{U,\iota}h}}{\mathbb{E}^{{\mathbb{Q}}_{0}}\left[e^{-\frac{1}{2}X_{0}^{\prime}P_{U,\iota}X_{0}+X_{0}^{\prime}P_{U,\iota}h}\right]}\right)\bigg{|}_{h=H_{0,\iota}}.$
Using this in (54) and that $X_{0}\stackrel{{\scriptstyle{\mathbb{Q}}_{0}}}{{\sim}}N(0,P_{X}^{-1})$, on $\left\{H_{0,\iota}=h\right\}$ the uninformed trader minimizes over $\theta=\theta(h)$
(75) $\begin{split}&\frac{1}{2}\theta^{\prime}\left(P_{U,\iota}+P_{X})^{-1}\right)\theta-\theta^{\prime}(P_{U,\iota}+P_{X})^{-1}\left(P_{U,\iota}h-(P_{U,\iota}+P_{X})p_{\iota}(h)\right).\end{split}$
(55) is immediate. ∎
@@
Throughout we suppress the dependence of all functions on ${\mathcal{Y}}$. First, (58) follows directly from (57) after dividing by $\alpha_{U}$ and right-multiplying by ${\mathcal{M}}^{-1}$. As such, provided there is a solution $\widehat{{\mathcal{Y}}}$ the price impact equilibrium follows with market signal $H_{\iota}$ from (59) and price $p_{\iota}$ from (60). The remainder of the proof identifies the policies in (61) starting with the uninformed trader. Here, using $\widehat{\theta}_{U,\iota}$ from (55), (47) and $\widehat{\psi}_{U,\iota}=\widehat{\Pi}+\widehat{\theta}_{U,\iota}$ we obtain
$\begin{split}\widehat{\psi}_{U,\iota}&=\widehat{\Pi}+P_{U,\iota}\left(H_{\iota}-p_{{\mathbb{Q}}_{0}}\right)-\left(P_{U,\iota}+P_{X}\right)\left(p_{\iota}(H_{\iota})-p_{{\mathbb{Q}}_{0}}\right),\\ &=\widehat{\Pi}+P_{X}p_{{\mathbb{Q}}_{0}}+P_{U,\iota}H_{\iota}-\left(P_{U,\iota}+P_{X}\right)p_{\iota}(H_{\iota}).\end{split}$
As for $I$ our goal is to write the strategy in the same form as for $U$, i.e. as a linear function of $\widehat{\Pi},p_{{\mathbb{Q}}_{0}},G_{0}$ and $p_{\iota}$. Starting with $\widehat{\theta}_{I,\iota}$ in (73) (at $g=G_{0}=G-p_{{\mathbb{Q}}_{0}}$ and using that in equilibrium $V_{0,p,\iota}=0$) and $\widehat{\psi}_{I,\iota}=\widehat{\Pi}+\widehat{\theta}_{I,\iota}$ we find
$\widehat{\psi}_{I,\iota}=\widehat{\Pi}+P_{X|G}^{1/2}(\widehat{{\mathcal{Y}}}+\widehat{{\mathcal{Y}}}^{\prime}+1_{d})^{-1}P_{X|G}^{-1/2}\bigg{(}P_{I}(G-p_{{\mathbb{Q}}_{0}})-P_{X|G}^{1/2}\widehat{{\mathcal{Y}}}P_{X|G}^{-1/2}\frac{Z_{N}}{\alpha_{I}}\bigg{)}$
Next, from (47), (52) and (53) at $V_{0,p,\iota}=0$ we find
@@
$\begin{split}Z_{N}&=\Lambda_{\iota}^{-1}{\mathcal{M}}^{-1}M_{p,\iota}^{-1}\left(p_{\iota}(H_{\iota})-p_{{\mathbb{Q}}_{0}}\right)-\Lambda_{\iota}^{-1}\left(G-p_{{\mathbb{Q}}_{0}}\right).\end{split}$
This gives $\widehat{\psi}_{I,\iota}=\widehat{\Pi}+\mathbf{M}_{1}p_{{\mathbb{Q}}_{0}}+\mathbf{M}_{2}G+\mathbf{M}_{3}p_{\iota}(H_{\iota})$ where
@@
$\begin{split}\mathbf{M}_{1}=P_{X|G}^{1/2}(\widehat{{\mathcal{Y}}}+\widehat{{\mathcal{Y}}}^{\prime}+1_{d})^{-1}P_{X|G}^{-1/2}\left(-P_{I}-P_{X|G}^{1/2}\widehat{{\mathcal{Y}}}P_{X|G}^{-1/2}\frac{1}{\alpha_{I}}\Lambda_{\iota}^{-1}\left(1_{d}-{\mathcal{M}}^{-1}M_{p,\iota}^{-1}\right)\right).\end{split}$
From (49), (51) we find
@@
$\begin{split}\mathbf{M}_{1}&=P_{X|G}^{1/2}(\widehat{{\mathcal{Y}}}+\widehat{{\mathcal{Y}}}^{\prime}+1_{d})^{-1}P_{X|G}^{-1/2}\bigg{(}-P_{I}-P_{X|G}^{1/2}\widehat{{\mathcal{Y}}}(1_{d}+\widehat{{\mathcal{Y}}}^{\prime})^{-1}P_{X|G}^{-1/2}P_{I}\bigg{(}1_{d}\\ &\qquad\qquad-P_{I}^{-1}P_{X|G}^{1/2}(1_{d}+\widehat{{\mathcal{Y}}}+\widehat{{\mathcal{Y}}}^{\prime})\widehat{{\mathcal{Y}}}^{-1}P_{X|G}^{1/2}\bigg{)}\bigg{)},\\ &=P_{X|G}^{1/2}(\widehat{{\mathcal{Y}}}+\widehat{{\mathcal{Y}}}^{\prime}+1_{d})^{-1}P_{X|G}^{-1/2}\bigg{(}-P_{X|G}^{1/2}(1_{d}+\widehat{{\mathcal{Y}}}+\widehat{{\mathcal{Y}}}^{\prime})(1_{d}+\widehat{{\mathcal{Y}}}^{\prime})^{-1}P_{X|G}^{-1/2}P_{I}\\ &\qquad\qquad+P_{X|G}^{1/2}(1_{d}+\widehat{{\mathcal{Y}}}+\widehat{{\mathcal{Y}}}^{\prime})(1_{d}+\widehat{{\mathcal{Y}}}^{\prime})^{-1}P_{X|G}^{1/2}\bigg{)},\\ &=P_{X|G}^{1/2}(1_{d}+\widehat{{\mathcal{Y}}}^{\prime})^{-1}P_{X|G}^{-1/2}(-P_{I}+P_{X|G})=P_{X|G}^{1/2}(1_{d}+\widehat{{\mathcal{Y}}}^{\prime})^{-1}P_{X|G}^{-1/2}P_{X}.\end{split}$
Next, we have
$\begin{split}\mathbf{M}_{2}&=P_{X|G}^{1/2}(\widehat{{\mathcal{Y}}}+\widehat{{\mathcal{Y}}}^{\prime}+1_{d})^{-1}P_{X|G}^{-1/2}\bigg{(}P_{I}+P_{X|G}^{1/2}\widehat{{\mathcal{Y}}}P_{X|G}^{-1/2}\frac{1}{\alpha_{I}}\Lambda_{\iota}^{-1}\bigg{)},\\ &=P_{X|G}^{1/2}(\widehat{{\mathcal{Y}}}+\widehat{{\mathcal{Y}}}^{\prime}+1_{d})^{-1}P_{X|G}^{-1/2}\bigg{(}P_{I}\alpha_{I}\Lambda_{\iota}+P_{X|G}^{1/2}\widehat{{\mathcal{Y}}}P_{X|G}^{-1/2}\bigg{)}\frac{1}{\alpha_{I}}\Lambda_{\iota}^{-1},\\ &=P_{X|G}^{1/2}(\widehat{{\mathcal{Y}}}+\widehat{{\mathcal{Y}}}^{\prime}+1_{d})^{-1}P_{X|G}^{-1/2}\bigg{(}P_{X|G}^{1/2}(1_{d}+\widehat{{\mathcal{Y}}}^{\prime})P_{X|G}^{-1/2}+P_{X|G}^{1/2}\widehat{{\mathcal{Y}}}P_{X|G}^{-1/2}\bigg{)}\frac{1}{\alpha_{I}}\Lambda_{\iota}^{-1},\\ &=\frac{1}{\alpha_{I}}\Lambda_{\iota}^{-1}.\end{split}$
Lastly, we have
$\begin{split}\mathbf{M}_{3}&=-P_{X|G}^{1/2}(\widehat{{\mathcal{Y}}}+\widehat{{\mathcal{Y}}}^{\prime}+1_{d})^{-1}P_{X|G}^{-1/2}\times P_{X|G}^{1/2}\widehat{{\mathcal{Y}}}P_{X|G}^{-1/2}\frac{1}{\alpha_{I}}\Lambda_{\iota}^{-1}{\mathcal{M}}^{-1}M_{p,\iota}^{-1},\\ &=-P_{X|G}^{1/2}(\widehat{{\mathcal{Y}}}+\widehat{{\mathcal{Y}}}^{\prime}+1_{d})^{-1}\widehat{{\mathcal{Y}}}(1_{d}+\widehat{{\mathcal{Y}}}^{\prime})^{-1}(1_{d}+\widehat{{\mathcal{Y}}}+\widehat{{\mathcal{Y}}}^{\prime})\widehat{{\mathcal{Y}}}^{-1}P_{X|G}^{1/2},\\ &=-P_{X|G}^{1/2}(1_{d}+\widehat{{\mathcal{Y}}}^{\prime})^{-1}P_{X|G}^{1/2}.\end{split}$
@@
Assume $H_{0}$ takes the form $H_{0}=G_{0}+\Lambda Z_{N}$ for some to-be-determined matrix $\Lambda$ and that the price is $p_{0}=p_{0}(H_{0})$. Then $\sigma(G_{0},H_{0})=\sigma(G_{0},Z_{N})$, and using the ${\mathbb{Q}}_{0}$ conditional law of $X_{0}$ given $(G_{0},Z_{N})$ in (71) as well as $X_{0}\stackrel{{\scriptstyle{\mathbb{Q}}_{0}}}{{\sim}}N(0,P_{X}^{-1})$, for the insider the problem in (64) is equivalent to minimizing over $\theta$
(76) $\begin{split}&\frac{1}{2}\theta^{\prime}(P_{I}+P_{X})^{-1}\theta-\theta^{\prime}(P_{I}+P_{X})^{-1}\left(P_{I}G_{0}-(P_{I}+P_{X})p_{0}(H_{0})\right).\end{split}$
The optimizer is $\widehat{\theta}_{I}=P_{I}G_{0}-(P_{I}+P_{X})p_{0}(H_{0})$. Similarly, as the uninformed signal is $H_{0}=X_{0}+(Y_{I}+\Lambda Z_{N})$ it is of the same form as the insiders, just with precision $P_{U}$ instead of $P_{I}$. As such, the uninformed trader’s optimal policy is $\widehat{\theta}_{U}=P_{U}H_{0}-(P_{U}+P_{X})p_{0}(H_{0})$. The clearing condition (44) is
$0=\alpha_{I}P_{I}\left(G_{0}+\frac{1}{\alpha_{I}}P_{I}^{-1}Z_{N}\right)+\alpha_{U}P_{U}H_{0}-(\alpha_{I}(P_{I}+P_{X})+\alpha_{U}(P_{U}+P_{X}))p_{0}(H_{0}).$
This gives $\Lambda=(1/\alpha_{I})P_{I}^{-1}$ and hence the market signal in (65) as well as precision $p_{U}$ in (66). $p_{0}$ takes the form
$p_{0}(H_{0})=(\alpha_{I}(P_{I}+P_{X})+\alpha_{U}(P_{U}+P_{X}))^{-1}\big{(}\alpha_{I}P_{I}+\alpha_{U}P_{U})H_{0},$
which yields the price in (67). Lastly, using $\widehat{\psi}_{I}=\widehat{\theta}_{I}+\widehat{\Pi}$ and $\widehat{\psi}_{U}=\widehat{\theta}_{U}+\widehat{\Pi}$ and $p(H)=p_{o}(H_{0})-p_{{\mathbb{Q}}_{0}}$ we obtain (68). ∎
The proof of Proposition A.7, as well as those in both Sections B.3 and B.4 simplify if we adjust the notation in (7) by defining
(77) $\beta:=\kappa p_{I}=\alpha_{I}^{2}p_{I}p_{N};\hskip 5.0pt\lambda_{I}:=\frac{\alpha_{I}}{\alpha_{I}+\alpha_{U}};\hskip 5.0pt\lambda_{U}:=\frac{\alpha_{U}}{\alpha_{I}+\alpha_{U}},\hskip 5.0ptR:=\frac{1}{1+p_{I}}.$
@@
Next, under Assumption A.4 the constants $p_{U,\iota}$ from (63) and $p_{U}$ from (69) are
(79) $\begin{split}p_{U,\iota}:=\frac{(1-R)\beta}{R(\beta+(1+\widehat{y})^{2})};\qquad p_{U}=\frac{(1-R)\beta}{R(1+\beta)}.\end{split}$
The matrices of Proposition A.6 are
(80) $\begin{split}\alpha_{I}P_{I}+\alpha_{U}P_{U}&=\frac{(\alpha_{I}+\alpha_{U})(1-R)(\lambda_{I}+\beta)}{R(1+\beta)}P_{X},\\ \alpha_{I}(P_{I}+P_{X})+\alpha_{U}(P_{U}+P_{X})&=\frac{(\alpha_{I}+\alpha_{U})(\lambda_{I}+\beta+R\lambda_{U})}{R(1+\beta)}P_{X}.\end{split}$
Lastly, from (39), (47), Assumption A.4 and (77) we obtain
(81) $X_{0}=P_{X}^{-1/2}\mathcal{E}_{X};\quad G_{0}=X_{0}+\sqrt{\frac{R}{1-R}}P_{X}^{-1/2}\mathcal{E}_{I};\qquad\frac{Z_{N}}{\alpha_{I}}=\sqrt{\frac{1-R}{\beta R}}P_{X}^{1/2}\mathcal{E}_{N},$
where $\mathcal{E}_{X},\mathcal{E}_{I},\mathcal{E}_{N}$ are three independent $N(0,1_{d})$ random variables under ${\mathbb{Q}}_{0}$. This implies we can write $H_{0}$ from (52) as (see (79))
(82) $H_{0}=X_{0}+\sqrt{\frac{R}{1-R}}P_{X}^{-1/2}\mathcal{E}_{I}+\frac{R}{\alpha_{I}(1-R)}P_{X}^{-1}Z_{N}.$
@@
We use the notation in (77), and suppress $\widehat{y}$ from $M_{p,\iota},{\mathcal{M}}$. We start with the price taking equilibria. Using (80) and (82) the price $p(H)$ from (67) is